2) Suppose that consumers of a good can be represented by the demand function Q(P)=

50-P. The good is manufactured by an upstream monopolist with cost function C(Q)=

Q2 + 2Q + 10. A downstream monopolist resells the good to consumers (without further

production activity).

(a) Determine the industry outcome, profits and consumer surplus.

(b) Consider a vertical merger. Compare the industry outcome, profits and consumer

surplus to part (a).

(c) Suppose the upstream monopolist franchises the product to the downstream firm.

Which two-part tariff should the upstream monopolist choose? Determine the profits of the

firms.


1
Expert's answer
2021-08-20T08:49:18-0400

"p=50-Q"

"TR=pQ"

"TR=50Q-Q^2"

"MR=50-2Q"

"MC=2Q+2"

"50-2Q=2Q+2"

"Q=12"

"p=38"

"Profit=278"

"CS=\frac{(50-38)\times12}{2}=72"